OpenStudy (anonymous):

name the property of real numbers illustrated by the equation pi times 3 = 3 times pie

OpenStudy (anonymous):

mmm pie

OpenStudy (anonymous):

\[\pi \times 3=3\times \pi\] the symbols change places, it is the "commutative law" for multiplication

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

oh, they call it "property" now

OpenStudy (anonymous):

btw the greek letter is spelled "pi" don't let your teacher see you writing "pie" he/she will think you are missing something

OpenStudy (anonymous):

commutative property of multiplication

OpenStudy (anonymous):

yes, that one

OpenStudy (anonymous):

ow thanks again